OpenStudy (anonymous):

add or subtract as indicated and write the result in standard form (2-7i)+(6+5i)+(2+5i)

OpenStudy (anonymous):

You mean like this 10+3i?

OpenStudy (anonymous):

i dont know i dont understand it

OpenStudy (anonymous):

Ya that's right, I will try to explain it

OpenStudy (anonymous):

To add two complex numbers, add their real parts and add their imaginary parts: (a 1 + b 1 i) + (a 2 + b 2 i) = (a 1 + a 2) + (b 1 + b 2)i .

OpenStudy (anonymous):

Don't forget to click on the best response button :)

OpenStudy (anonymous):

and i need the work i dont know how to do it i even had my sister whose in college try to help me but she didnt understand.

OpenStudy (anonymous):

The formula I posted shows you exactly how you need to work it out to get the answer

OpenStudy (anonymous):

Here is an example using the formula i posted (12 + 6i) + (11 + 5i) = (12 + 11) + (6 + 5)i = 23 + 11i

OpenStudy (anonymous):

I know you have three equations so just add an extra variable a3 and b3i